# Atomically clean free-standing two-dimensional materials through heating in ultra-high vacuum Supporting Information for Atomically clean free-standing two-dimensional materials through heating in ultra-high vacuum

### v1

@@ Methods
The initially contaminated sample, where we measured an average clean area of only around 10% (Figure 3a), was annealed at 200 °C (Figure 3b), after which around 30% of the free-standing area was identified to be atomically clean monolayer h-BN. An example of atomically clean and defect-free h-BN annealed at 200 °C is shown in Figure 3c. In addition to contaminants adhered to the step edges between monolayer and bilayer regions, large portions of the surface are covered by networks of carbonaceous contamination with embedded clean areas in the range of only hundreds of nm2. These networks appeared to get more and more fragmented as the temperature was increased to 250 °C and 300 °C without substantially increasing the total clean area. Only when raising the temperature to 400 °C or higher did these networks disappear across all imaged suspended monolayer areas in the sample, and the fraction of the clean suspended monolayer area approached 50%. We note that, accounting for the estimated relative free-standing monolayer area of ca. 50%, the monolayer areas exhibit a cleanliness approaching 100%, similar to what we found for graphene transferred via the same transfer method, annealed at the same temperatures. Similar to graphene, the residual observed contamination takes the form of nanosized islands, often but not always, around copper and iron nanoclusters (Supplementary Figure 12). Also, we identify many small carbon-rich contamination agglomerations devoid of metallic residues within otherwise clean areas, which are likely attached to point defects in the material. However, the much higher density of these islands compared to what we found in graphene made via the same transfer method suggests a higher defect density of our transferred h-BN. While no systematic study was carried out here to characterize the defect density before or after the annealing steps, we did not observe a higher defect density in the atomically clean sites after annealing than before the treatment (Supplementary Figure 16, Supporting Information), suggesting that UHV annealing can remove the majority of the surface contamination without harming the structural integrity of free-standing monolayer h-BN in the process. The higher defect density toward the center of the frames in Supplementary Figure 16 (Supporting Information) originates from electron-beam-induced damage[84] from prior imaging at higher local current densities to acquire the images in the insets. While it was easy to find atomically clean areas that were 200x200 nm2 in size for graphene after the 200 °C bake, we have a higher amount of residual contamination on h-BN at the same temperature compared to graphene. The differences in the amount and morphology of contamination on h-BN and graphene after baking at only 200 °C further support a higher defect density of h-BN. While we did not investigate whether this is due to a lower quality of the source material or if the defects were introduced during sample transfer, recent literature also reported on a relatively high defect density (ca. ) in monolayer h-BN transferred via electrochemical delamination.[74] Given the high thermal stability of h-BN,[85, 86] and the lack of reactive species in our annealing atmosphere, in conjunction with only a modest annealing temperature, we presume that the 200 °C bake was not the source of our observed defects.
@@
To get a better understanding of the removal of only airborne (not polymer-induced) contamination, a cleaned graphene sample that was previously heated to 450 °C (Figure 4a) was removed from the UHV system for 15 min, during which it was exposed to ambient laboratory air, but not to polymers. After re-introducing it into UHV through the regular ca. 160 °C vacuum bake overnight, the sample was found to again be heavily contaminated, with practically all of the surface being covered by hydrocarbon species (Figure 4b). However, Figure 4c shows that re-heating the sample to 200 °C already revealed more clean areas than the previous 200 °C bake (Figure 1c), where the contamination also consisted of PMMA residue. Then, the temperature was iteratively increased to higher values. While raising the temperature to 250 °C (Figure 4d) did not increase the clean area, and 300 °C (Figure 4e) provided only modest improvements, uncovering more and more of the pristine graphene lattice, it was only at 350 °C (Figure 4f) that the sample exhibited the same level of cleanliness as after the previous 450 °C bake. This can be explained by the lack of required depolymerization of the PMMA, which is initiated by the breaking of the backbone of the PMMA, reportedly occurring between 350 °C and 400 °C,[87, 42, 27, 49] as the polymer residue had been removed by the previous 450 °C bake (before removal from UHV), with mostly airborne hydrocarbons on the surface after the re-contamination step.
@@ Conclusions
We report that thermal annealing in UHV is an effective method to remove airborne hydrocarbon and polymer contamination from both graphene and h-BN surfaces without compromising the integrity of the materials. Even temperatures as low as 200 °C can lead to noticeable contamination removal, especially when there is no polymer or metallic residue present, where well over 50% of the probed free-standing monolayer area is identified as atomically clean. This lower annealing temperature can be a viable cleaning step for fundamental study and manipulation of 2D materials, including impurity atom implantation and defect-engineering, where heating the specimen to higher temperatures might not be necessary, and may only amplify the effect of the mismatch of the thermal expansion coefficient of the material and sample support, resulting in excess strain, induced corrugations, and tearing of the material. However, heating to higher temperatures is needed to achieve cleaner surfaces as the cleaning mechanism is incomplete until the materials are annealed at over 400 °C, where the decomposition and subsequent desorption of PMMA introduced during the sample transfer has occurred, and almost all of the contamination is removed. There, we find that achieving a higher degree of cleanliness is predominantly limited by defects in the material, such as grain boundaries and point defects, as well as residual metal clusters introduced during the sample growth and transfer process. A high amount of metallic residue can be detrimental to the achievable cleanliness, as we found that a sample mostly devoid of metal clusters when annealed at only 200 °C can exhibit similar levels of cleanliness as a heavily metal-contaminated one that is baked at over 450 °C. Hence, we suggest to avoid using metal-based etchants during the sample transfer without further steps to dissolve the metallic residue, which UHV annealing cannot remove. Also, when there are no polymer residues on the surface, we obtain cleaner surfaces at lower annealing temperatures, where heating graphene to 350 °C was sufficient in obtaining atomically clean areas thousands of nm2 in size, compared to over 400 °C if there was polymer present.
@@
We also conclude that h-BN exhibits a thermal response to surface cleaning in UHV that is comparable to graphene. At temperatures above 400 °C, the surfaces are almost fully devoid of contamination, with contamination anchored to defect sites and metal clusters being the main limitation, similar to graphene. However, the quality of the material used in this work was found to be vastly inferior to that of graphene when samples transferred using the same transfer method were compared. Not only does the high abundance of multilayer structures originating from the growth heavily limit the amount of monolayer h-BN, but we also presume that the smaller average grain sizes of h-BN compared to graphene cause more damage to the material during the transfer. The resulting multilayer structures from the material folding over, covering up what would otherwise have been pristine monolayer regions, trap contaminants between the layers, making them difficult, if not impossible, to remove.
@@
Finally, we observe only a minimal to no increase in defects for both graphene and h-BN after annealing, further reinforcing that UHV annealing can reliably clean surfaces without unintended defect creation. These large, clean surfaces can be beneficial for applications requiring large pristine areas, or for subsequent nanoscale engineering steps like defect-engineering and doping, all of which benefit from large exposed clean areas to not only increase the likelihood of interacting with the material rather than contamination, but to also utilize the improved physical properties of contamination-free 2D materials.
@@ Acknowledgments
P.I., D.L., and C.M. designed the UHV heating chamber, with further contribution from C.S. and M.L. when integrating it into the UHV system. S.C. and P.I. developed and integrated the sample transfer process, with P.I. also preparing the samples. T.A.B. prepared further hexagonal boron nitride samples used for testing the UHV heating stage. P.I. carried out the imaging for all samples. P.I. and D.L. conceptualized the data analysis methodology, with P.I. conducting the data analysis. J.K. supervised the study. The manuscript text was initially written by P.I., and improved upon by all authors.
@@
The experimental data supporting these findings will be openly available at PHAIDRA at https://phaidra.univie.ac.at/o:2149991 upon publication.[88]
@@
See Supporting Information for additional discussion regarding mobile contamination, sample support comparison, applied false color grading of ADF-STEM images, and a thorough description of the digital image processing steps. Additional figures on the MAADF-STEM image acquisition process, EEL spectra of metal and SiO2 contaminants, h-BN supplier comparison and defect observations, and design sketches of the UHV heating chamber are also provided.
@@
Philipp Irschik1,∗, David Lamprecht1,2, Shrirang Chokappa1, Clemens Mangler1, Carsten Speckmann1, Thuy An Bui1, Manuel Längle1, Lado Filipovic2, Jani Kotakoski1,∗
1Faculty of Physics, University of Vienna, Boltzmanngasse 5,
1090 Vienna, Austria
2Institute for Microelectronics, TU Wien, Gußhausstraße 25-29, 1040 Vienna, Austria
@@
At higher annealing temperatures, we observe substantial warping and damage of the structural sample support, likely occurring from the mismatch of the thermal expansion coefficient of the gold bars and amorphous carbon film, in addition to thinning of the already thin (ca. 10 nm) amorphous carbon support membrane. Figure 9a shows the observed compressions of the imaged projections of the holes in the support film, which manifest as elliptic rather than circular appearing holes, originating from thermally-induced stretching of the sample support. This not only causes mechanical strain, creating tears and cracks in the graphene (red arrows in Figure 9a), but also makes it more difficult to get larger areas into focus at once with the scanning probe due to strong height differences within each hole (Figure 9b), in addition to support-induced vibrations which manifest in non-ideal imaging conditions, limiting acquisition of atomically-resolved images. Therefore, another graphene sample prepared via electrochemical delamination was transferred onto a holey silicon nitride support membrane with a rigid 200 nm membrane thickness and was heated at 450 °C for 2 h in UHV. After annealing at 450 °C, the sample support was still in pristine condition (Figure 9c), mitigating support-induced corrugations of graphene (Figure 9d) and thus also minimizing strain-induced damage to the material. This sample exhibited a similar level of cleanliness as the one on the holey carbon support film that had also been annealed at 450 °C, with a measured average clean area above 95% (Figure 1a, main text), slightly higher than the 90% measured from the other sample support. However, we attribute this minimal difference to local differences from the source material and transfer, in addition to contamination overestimation arising from unfocused features caused by the earlier-mentioned height variations in the other samples, rather than due to a different thermal response of the Si/SiNx compared to the holey amorphous carbon support.

### v2

@@ Methods
The as-prepared sample, where we measured an average clean area of only around 10% (Figure 3a), was annealed at 200 °C (Figure 3b), after which around 30% of the free-standing area was identified to be atomically clean monolayer h-BN. An example of atomically clean and defect-free h-BN annealed at 200 °C is shown in Figure 3c. In addition to contaminants adhered to the step edges between monolayer and bilayer regions, large portions of the surface are covered by networks of carbonaceous contamination with embedded clean areas in the range of only hundreds of nm2. These networks appeared to become more and more fragmented as the temperature was increased to 250 °C and 300 °C without substantially increasing the total clean area. Only when raising the temperature to 400 °C or higher did these networks disappear across all imaged suspended monolayer areas in the sample, and the fraction of the clean suspended monolayer area approached 50%. We note that, accounting for the estimated relative free-standing monolayer area of ca. 50%, the monolayer areas exhibit a cleanliness approaching 100%, similar to what we found for graphene transferred via the same transfer method, annealed at the same temperatures. Similar to graphene, the residual observed contamination takes the form of nanosized islands, often but not always, around copper and iron nanoclusters (Supplementary Figure 12). We also identify many small carbon-rich contamination agglomerations devoid of metallic residues within otherwise clean areas, which are likely attached to point defects in the material. However, the much higher density of these islands compared to what we found in graphene made via the same transfer method suggests a higher defect density of our transferred h-BN. While no systematic study was carried out here to characterize the defect density before or after the annealing steps, we did not observe a higher defect density in the atomically clean sites after annealing than before the treatment (Supplementary Figure 16, Supporting Information), suggesting that UHV annealing can remove the majority of the surface contamination without harming the structural integrity of free-standing monolayer h-BN in the process. The higher defect density toward the center of the frames in Supplementary Figure 16 (Supporting Information) originates from electron-beam-induced damage[85] from prior imaging at higher local current densities to acquire the images in the insets. While it was straightforward to find atomically clean areas which were 200x200 nm2 in size for graphene after the 200 °C bake, we observe a higher amount of residual contamination on h-BN that was annealed at the same temperature. The differences in the amount and morphology of contamination on h-BN and graphene after baking at only 200 °C further support a higher defect density of h-BN. While we did not investigate whether this is due to a lower quality of the source material or if the defects were introduced during sample transfer, recent literature also reported on a relatively high defect density (ca. ) in monolayer h-BN transferred via electrochemical delamination[75]. Given the high thermal stability of h-BN [86, 87], and the lack of reactive species in our annealing atmosphere, in conjunction with only a modest annealing temperature, we presume that the 200 °C bake was not the source of our observed defects.
@@
To get a better understanding of the removal of only airborne (not polymer-induced) contamination, a cleaned graphene sample that was previously heated to 450 °C (Figure 4a) was removed from the UHV system for 15 min, during which it was exposed to ambient laboratory air, but not to polymers. After re-introducing it into UHV through the regular ca. 160 °C vacuum bake overnight, the sample was found to again be heavily contaminated, with practically all of the surface being covered by hydrocarbon species (Figure 4b). However, Figure 4c shows that re-heating the sample to 200 °C already revealed more clean areas than the previous 200 °C bake (Figure 1c), where the contamination also consisted of PMMA residue. Then, the temperature was iteratively increased to higher values. While raising the temperature to 250 °C (Figure 4d) did not increase the clean area, and 300 °C (Figure 4e) provided only modest improvements, uncovering more and more of the pristine graphene lattice, it was only at 350 °C (Figure 4f) that the sample exhibited the same level of cleanliness as after the previous 450 °C bake. This can be explained by the lack of required depolymerization of the PMMA, which is initiated by the breaking of the backbone of the PMMA, reportedly occurring between 350 °C and 400 °C[88, 42, 27, 49], as the polymer residue had been removed by the previous 450 °C bake (before removal from UHV), with mostly airborne hydrocarbons on the surface after the re-contamination step.
@@ Conclusions
We report that thermal annealing in UHV is an effective method to remove airborne hydrocarbon and polymer contamination from both graphene and h-BN surfaces without compromising the integrity of the materials. Even annealing at temperatures as low as 200 °C can lead to noticeable contamination removal, especially when there is no polymer or metallic residue present, where well over 50% of the probed free-standing monolayer surface is identified as atomically clean. This lower annealing temperature can be a viable cleaning step for fundamental study and manipulation of 2D materials, including impurity atom implantation and defect-engineering, where heating the specimen to higher temperatures might not be necessary or possible, and may only amplify the effect of the mismatch of the thermal expansion coefficient of the material and sample support, resulting in excess strain, induced corrugations, and tearing of the material. However, heating to higher temperatures is needed to achieve cleaner surfaces as the cleaning mechanism is incomplete until the materials are annealed at over 400 °C, where the decomposition and subsequent desorption of PMMA introduced during the sample transfer has occurred, and almost all of the contamination is removed. There, we find that achieving a higher degree of cleanliness is predominantly limited by defects in the material, such as grain boundaries and point defects, as well as residual metal clusters introduced during the sample growth and transfer processes. A high amount of metallic residue can be detrimental to the achievable cleanliness, as we found that a sample mostly devoid of metal clusters, when annealed at only 200 °C, can exhibit similar levels of cleanliness as a heavily metal-contaminated one that is baked at over 450 °C. Hence, we suggest to avoid using metal-based etchants during the sample transfer without further steps to dissolve the metallic residue, which UHV annealing cannot remove. Also, when there are no polymer residues on the surface, we obtain cleaner surfaces at lower annealing temperatures, where heating graphene to 350 °C was sufficient in obtaining atomically clean areas thousands of nm2 in size, compared to over 400 °C if there was polymer present.
@@
We also conclude that h-BN exhibits a thermal response to surface cleaning in UHV that is comparable to graphene. At temperatures above 400 °C, the surfaces are almost fully devoid of contamination, with contamination anchored to defect sites and metal clusters being the main limitation, similar to graphene. However, the quality of the material used in this work was found to be vastly inferior to that of graphene when samples transferred using the same transfer method were compared. Not only does the high abundance of multilayer structures originating from the growth heavily limit the amount of monolayer h-BN, but we also presume that the smaller average grain sizes of h-BN compared to graphene cause more damage to the material during the transfer. The resulting multilayer structures from the material folding over, covering up what would otherwise have been pristine monolayer regions, and inhibit contamination removal.
@@
Finally, we observe only a minimal to no increase in defects for both graphene and h-BN after annealing, further reinforcing that UHV annealing can reliably clean surfaces without unintended defect creation. These large, clean surfaces can be beneficial for applications requiring large pristine areas, or for subsequent nanoscale engineering steps like defect-engineering and doping, all of which benefit from large exposed clean areas to not only increase the likelihood of interacting with the material rather than contamination, but to also utilize the improved physical properties of contamination-free 2D materials. Extending this cleaning procedure to other classes of 2D materials with different defect tolerances and less thermal stability, such as transition metal dichalcogenides[89, 90, 91], is an important next step in evaluating its applicability for a broader range of materials, as we find that temperatures above 400 °C are necessary to remove polymer contamination from graphene and h-BN, resulting in nearly residue-free, pristine surfaces.
@@ Acknowledgments
P.I., D.L., and C.M. designed the UHV heating chamber, with further contribution from C.S. and M.L. when integrating it into the UHV system. S.C. and P.I. developed and integrated the sample transfer process, with P.I. also preparing the samples. T.A.B. prepared further hexagonal boron nitride samples used for testing the UHV heating stage. P.I. carried out the imaging for all samples. P.I. and D.L. conceptualized the data analysis methodology, with P.I. conducting the data analysis. J.K. and L.F. supervised the study. The manuscript text was initially written by P.I., and improved upon by all authors.
@@
The experimental data supporting these findings will be openly available at PHAIDRA at https://phaidra.univie.ac.at/o:2149991 upon publication[92].
@@
See Supporting Information for additional discussion regarding electron-beam-induced hydrocarbon deposition, sample support comparison, applied false color grading of ADF-STEM images, and a thorough description of the digital image processing steps. Additional figures on the MAADF-STEM image acquisition process, EEL spectra of metal and SiO2 contaminants, h-BN supplier comparison and defect observations, and design sketches of the UHV heating chamber are also provided.
@@
Philipp Irschik1,2,∗, David Lamprecht1,3, Shrirang Chokappa1,2, Clemens Mangler1, Carsten Speckmann1, Thuy An Bui1,2, Manuel Längle1, Lado Filipovic3, Jani Kotakoski1,∗
1University of Vienna, Faculty of Physics, Boltzmanngasse 5, 1090 Vienna, Austria
2University of Vienna, Vienna Doctoral School in Physics, Boltzmanngasse 5, 1090 Vienna, Austria
3Institute for Microelectronics, TU Wien, Gußhausstraße 27-29/E360, 1040 Vienna, Austria
@@
At higher annealing temperatures, we observe substantial warping and damage of the structural sample support, likely occurring from the mismatch of the thermal expansion coefficient of the gold bars and amorphous carbon film, in addition to thinning of the already thin (ca. 10 nm) amorphous carbon support membrane. Figure 9a shows the observed compressions of the imaged projections of the holes in the support film, which manifest as elliptic rather than circular appearing holes, originating from thermally-induced stretching of the sample support. This not only causes mechanical strain, creating tears and cracks in the graphene (red arrows in Figure 9a), but also makes it more difficult to get larger areas into focus at once with the scanning probe due to strong height differences within each hole (Figure 9b), in addition to support-induced vibrations which manifest in non-ideal imaging conditions, limiting acquisition of atomically-resolved images. Therefore, another graphene sample prepared via electrochemical delamination was transferred onto a holey silicon nitride support membrane with a rigid 200 nm membrane thickness and was heated at 450 °C for 2 h in UHV. After annealing at 450 °C, the sample support was still in pristine condition (Figure 9c), mitigating support-induced corrugations of graphene (Figure 9d) and thus also minimizing strain-induced damage to the material. This sample exhibited a similar level of cleanliness as the one on the holey carbon support film that had also been annealed at 450 °C, with a measured average clean area above 95% (Figure 1a, main text, purple dots), slightly higher than the 90% measured from the other sample support. However, we attribute this minimal difference to local differences from the source material and transfer, in addition to contamination overestimation arising from unfocused features caused by the earlier-mentioned height variations in the other samples, rather than due to a different thermal response of the Si/SiNx compared to the holey amorphous carbon support.